# Scalable Entanglement Detection in Quantum Systems via Fisher Linear Discriminant Analysis

### v1

@@ 1 Introduction
Machine learning has emerged as a powerful alternative for addressing complex quantum classification tasks, including entanglement detection [27, 28, 29, 30, 31, 32]. Supervised learning techniques, such as support vector machines (SVMs) and neural networks, have been successfully applied to distinguish entangled states from separable ones by learning from labeled training data [33, 34, 35, 36]. Unsupervised methods, including clustering and dimensionality reduction, have also been explored for identifying entanglement without prior state knowledge [37, 38, 39, 40]. However, many ML approaches face challenges in interpretability, computational efficiency, or generalization to unseen states [30, 41, 42, 18]. In this context, classical FLDA [43, 44] offers a compelling balance between simplicity and performance, providing a linear decision boundary that maximizes separability between classes while remaining computationally efficient.
@@ 2 FLDA
In discriminant analysis for classification, a new sample $\mathbf{x}\in\mathbb{R}^{n}$ is projected into a $k$-dimensional discriminant subspace defined by the matrix of discriminant vectors $\mathbf{W}\in\mathbb{R}^{n\times k}$. The projection is computed as $\mathbf{y}=\mathbf{W}^{T}\mathbf{x}$, where $\mathbf{y}\in\mathbb{R}^{k}$ represents the sample’s coordinates in the reduced-dimensional subspace. When $\mathbf{W}$ is orthonormal, satisfying $\mathbf{W}^{T}\mathbf{W}=\mathbf{I}_{k}$, this mapping is an orthogonal projection, ensuring that $\mathbf{W}\mathbf{y}$ is the point in the subspace closest to $\mathbf{x}$ in terms of Euclidean distance. For the special case where $k=1$, the matrix $\mathbf{W}$ reduces to a single vector $\mathbf{w}\in\mathbb{R}^{n}$, and the projection $y=\mathbf{w}^{T}\mathbf{x}$ yields a scalar, representing the coordinate of $\mathbf{x}$ along the line spanned by $\mathbf{w}$. This dimensionality reduction preserves or enhances class separability, facilitating the use of simplified decision boundaries. Classification is achieved by assigning the sample to the class whose projected mean, $\mathbf{W}^{T}\boldsymbol{\mu}_{i}$, is closest to $\mathbf{y}$ in the subspace, typically measured using Euclidean distance.
@@ 3 Quantum Entanglement Detection
Feature vectors arise from measurement outcomes like probabilities $p_{i}=\text{Tr}(\rho\Pi_{i})$ or expectation values $\langle O\rangle=\text{Tr}(\rho O)$. Dimensionality $n$ scales with measurements, potentially exponentially with qubits. FLDA reduces this to k1k-1, and $\mathbf{W}$ identifies key measurements (e.g., Pauli operators).
@@
  • Noise Handling: Noise inflates $S_{W}$, requiring filtering or robust statistics.
@@ 3.1 Two-Qubit
$\rho_{Wer}^{(2)}=\frac{1}{4}\sum_{i,j=0}^{3}c_{ij}(\sigma_{i}\otimes\sigma_{j}),\quad c_{ij}=\delta_{ij}\cdot\begin{cases}1&\text{if }i=0,\\ ps_{i}&\text{if }i=1,2,3.\end{cases}$ (6)
@@
The entanglement of the two-qubit Werner state $\rho_{\text{Wer}}^{(2)}$ is assessed using the Peres-Horodecki PPT criterion. The critical eigenvalue condition $\frac{1-3p}{4}<0$ implies $p>\frac{1}{3}$. Thus, the state is separable for $p\leq\frac{1}{3}$ (PPT) and entangled for $p>\frac{1}{3}$, where the partial transpose yields a negative eigenvalue.
@@
Considering a new Werner state with unknown $p$, the same observables are measured, the feature vector $x$ is constructed, and projected to $y=W^{T}x$. The state is classified as separable if $y$ is closer to $W^{T}\mu_{1}$, or entangled if closer to $W^{T}\mu_{2}$.
@@ 3.2 Three-Qubit Werner States
For three qubits, the classification of mixed three-qubit states, as detailed in [46], involves categorizing states based on their separability across different bipartitions using the PPT criterion. A three-qubit Werner state, based on a general entangled state $|\Psi_{GHZ}\rangle$, interpolates between a maximally entangled state and the maximally mixed state. In the Pauli basis, it is:
$\rho_{Wer}^{(3)}=\frac{1}{8}\sum_{i,j,k\in\{0,1,2,3\}}c_{ijk}(\sigma_{i}\otimes\sigma_{j}\otimes\sigma_{k}),c_{ijk}=\delta_{ij}\delta_{ik}\delta_{jk}\cdot\begin{cases}1&\text{if }i=0,\\ p\,s_{i}&\text{if }i=1,2,3.\end{cases}$ (9)

### v2

@@ 1 Introduction
Machine learning has emerged as a powerful alternative for addressing complex quantum classification tasks, including entanglement detection [27, 28, 29, 30, 31, 32]. Supervised learning techniques, such as support vector machines (SVMs) and neural networks, have been successfully applied to distinguish entangled states from separable ones by learning from labeled training data [33, 34, 35, 36]. Unsupervised methods, including clustering and dimensionality reduction, have also been explored for identifying entanglement without prior state knowledge [37, 38, 39, 40]. However, many ML approaches face challenges in interpretability, computational efficiency, or generalization to unseen states [30, 41, 42]. In this context, classical FLDA [43, 44] offers a compelling balance between simplicity and performance, providing a linear decision boundary that maximizes separability between classes while remaining computationally efficient.
@@ 2 FLDA
In discriminant analysis for classification, a new sample $\mathbf{x}\in\mathbb{R}^{n}$ is projected into a $k$-dimensional discriminant subspace defined by the matrix of discriminant vectors $\mathbf{W}\in\mathbb{R}^{n\times k}$. The projection is computed as $\mathbf{y}=\mathbf{W}^{T}\mathbf{x}$, where $\mathbf{y}\in\mathbb{R}^{k}$ represents the sample’s coordinates in the reduced-dimensional subspace which has at most k1k-1 dimensions. When $\mathbf{W}$ is orthonormal, satisfying $\mathbf{W}^{T}\mathbf{W}=\mathbf{I}_{k}$, this mapping is an orthogonal projection, ensuring that $\mathbf{W}\mathbf{y}$ is the point in the subspace closest to $\mathbf{x}$ in terms of Euclidean distance. For the special case where $k=1$, the matrix $\mathbf{W}$ reduces to a single vector $\mathbf{w}\in\mathbb{R}^{n}$, and the projection $y=\mathbf{w}^{T}\mathbf{x}$ yields a scalar, representing the coordinate of $\mathbf{x}$ along the line spanned by $\mathbf{w}$. This dimensionality reduction preserves or enhances class separability, facilitating the use of simplified decision boundaries. Classification is achieved by assigning the sample to the class whose projected mean, $\mathbf{W}^{T}\boldsymbol{\mu}_{i}$, is closest to $\mathbf{y}$ in the subspace, typically measured using Euclidean distance.
@@ 3 Quantum Entanglement Detection
Feature vectors arise from measurement outcomes like probabilities $p_{i}=\text{Tr}(\rho\Pi_{i})$ or expectation values $\langle O\rangle=\text{Tr}(\rho O)$.
The dimensionality $n$ of the full feature space (e.g., the number of all non-identity Pauli operators) scales exponentially with the number of qubits ($4^{N}-1$). FLDA reduces this to k1k-1, and $\mathbf{W}$ identifies key measurements (e.g., Pauli operators).
@@
  • Noise Handling: Noise typically increases the within-class scatter $S_{W}$ and can also distort the between-class scatter $S_{B}$, requiring filtering or robust statistics.
@@ 3.1 Two-Qubit
$c_{ij}=0$ for $i\neq j$, and for the diagonal terms $c_{ii}$ we have
$\rho_{Wer}^{(2)}=\frac{1}{4}\sum_{i,j=0}^{3}c_{ij}(\sigma_{i}\otimes\sigma_{j}),\quad c_{ii}=\begin{cases}1&\text{if }i=0,\\ ps_{i}&\text{if }i=1,2,3.\end{cases}$ (6)
@@
The entanglement of the two-qubit Werner state $\rho_{\text{Wer}}^{(2)}$ is assessed using the Peres-Horodecki criterion. The critical eigenvalue condition $\frac{1-3p}{4}<0$ implies $p>\frac{1}{3}$. Thus, the state is separable for $p\leq\frac{1}{3}$ and entangled for $p>\frac{1}{3}$, where the partial transpose yields a negative eigenvalue.
@@
$W$ defined in both Eq.7,Eq.8 is a $1\times 15$ projection vector. Considering a new Werner state with unknown $p$, the same observables are measured, the feature vector $x$ is constructed, and projected to $\mathbf{y}=\mathbf{W}^{T}\mathbf{x}$. The state is classified as separable if $\mathbf{y}$ is closer to $W^{T}\mu_{1}$, or entangled if closer to $W^{T}\mu_{2}$.
@@
Algorithm 1 Two-Qubit Entanglement Detection via FLDA
1:Quantum state parameters: $p$ (Werner) or $\theta_{0},\theta_{1}$ (concurrence)
2:Observables: $\{\sigma_{x},\sigma_{y},\sigma_{z},\sigma_{i}\otimes\sigma_{i}\}$ for $i=1,2,3$
[…]
4:Classification: “Entangled” or “Separable”
5:
6:function Detect Entanglement
7:  Prepare two-qubit quantum state
8:  if using Werner state then
9:   $\rho_{\text{Wer}}^{(2)}\leftarrow\frac{1}{4}\sum_{i,j=0}^{3}c_{ij}(\sigma_{i}\otimes\sigma_{j})$
10:   $c_{ij}\leftarrow\delta_{ij}\cdot\begin{cases}1&\text{if }i=0\\ ps_{i}&\text{if }i=1,2,3\end{cases}$
11:  else
12:   $|\psi\rangle\leftarrow\cos\left(\frac{\theta_{0}}{2}\right)|00\rangle-i\sin\left(\frac{\theta_{0}}{2}\right)\cos\left(\frac{\theta_{1}}{2}\right)|10\rangle-i\sin\left(\frac{\theta_{0}}{2}\right)\sin\left(\frac{\theta_{1}}{2}\right)|11\rangle$
13:  end if
14:  for each observable $O$ do
15:   Perform $N_{\text{trials}}$ measurements
16:   $\langle O\rangle\leftarrow\operatorname{tr}(\rho O)$ Estimate expectation value
17:  end for
18:  $\mathbf{x}\leftarrow[\langle O_{1}\rangle,\langle O_{2}\rangle,\dots,\langle O_{n}\rangle]^{T}$ Construct feature vector
19:  Normalize/standardize features
20:  if $p\leq\frac{1}{3}$ (Werner) or $C=0$ (concurrence) then
21:   $\text{label}\leftarrow\text{``Separable''}$
22:  else
23:   $\text{label}\leftarrow\text{``Entangled''}$
24:  end if
25:  Compute $S_{W}$ and $S_{B}$ matrices:
26:  $S_{B}\leftarrow\sum_{i=1}^{k}N_{i}(\boldsymbol{\mu}_{i}-\boldsymbol{\mu})(\boldsymbol{\mu}_{i}-\boldsymbol{\mu})^{T}$
27:  $S_{W}\leftarrow\sum_{i=1}^{k}\sum_{\mathbf{x}\in C_{i}}(\mathbf{x}-\boldsymbol{\mu}_{i})(\mathbf{x}-\boldsymbol{\mu}_{i})^{T}$
28:  Solve $S_{B}\mathbf{w}=\lambda S_{W}\mathbf{w}$ Generalized eigenvalue problem
29:  Obtain projection vector $\mathbf{w}$
30:  for new state $\mathbf{x}_{\text{new}}$ do
31:   $y\leftarrow\mathbf{w}^{T}\mathbf{x}_{\text{new}}$ Project onto discriminant direction
32:   if $y$ closer to $\mathbf{w}^{T}\boldsymbol{\mu}_{\text{entangled}}$ then
33:     return “Entangled”
34:   else
35:     return “Separable”
36:   end if
37:  end for
38:end function
@@ 3.2 Three-Qubit Werner States
For three qubits, the classification of mixed three-qubit states, as detailed in [46], involves categorizing states based on their separability across different bipartitions using the PPT criterion. A three-qubit Werner state, based on a general entangled state $|\Psi_{GHZ}\rangle$, interpolates between a maximally entangled state and the maximally mixed state. $c_{ijk}=0$ for $i\neq j\neq k$, and for the diagonal terms $c_{iii}$ we have
$\rho_{Wer}^{(3)}=\frac{1}{8}\sum_{i,j,k\in\{0,1,2,3\}}c_{ijk}(\sigma_{i}\otimes\sigma_{j}\otimes\sigma_{k}),c_{iii}=\begin{cases}1&\text{if }i=0,\\ p\,s_{i}&\text{if }i=1,2,3.\end{cases}$ (9)